# From Perceptions to Decisions: Wildfire Evacuation Decision Prediction with Behavioral Theory-informed LLMs

### v1

@@ Abstract
Ruxiao Chen*1, Chenguang Wang1thanks: Equal Contribution., Yuran Sun2, Xilei Zhao2, Susu Xu1 1Johns Hopkins University 2University of Florida {rchen117@jh.edu,cwang274, sxu83}@jh.edu
@@ 1 Introduction
Wildfires are emerging as a significant natural hazard worldwide Jain et al. (2020); Zahura et al. (2024). In the January 2025 Southern California wildfires, more than 200,000 residents received evacuation orders to leave their homes Stelloh et al. (2025). There is an urgent demand for emergency planners and policymakers to develop effective evacuation strategies to mitigate wildfire impacts Mockrin et al. (2018); Tapley et al. (2023). However, successful evacuations require a clear understanding of the human evacuation decision-making process and outcomes (i.e. whether individuals will follow the order to evacuate or stay) during these events to help policymakers improve evacuation order design, develop more efficient emergency response strategies, and build more resilient communities Collins et al. (2018); Lovreglio et al. (2020); Hong and Frias-Martinez (2020); Sun et al. (2024c).
@@
Previous studies often construct evacuation choice models through a conceptual framework, Protective Action Decision Model (PADM) Strahan and Watson (2019); Lovreglio et al. (2019); Santana et al. (2021); Sun et al. (2024a), which are designed to incorporate psychological factors, like individual risk perception and threat assessment, into the prediction process. Based on the PADM framework, past methods may employ various statistical models (e.g., logistic regression Forrister et al. (2024a), multinomial logistic regression McCaffrey et al. (2018)) to predict individual-evacuation decisions using socio-demographic information as inputs, trained on post-wildfire survey data. However, these traditional PADM-type statistical models lack reasoning capabilities to capture the diverse and complex logic underlying human decision-making due to limited data and restrictive modeling structure, even when the survey design is grounded in established behavioral theories. In addition, these statistical methods struggle to integrate qualitative descriptions — such as narrative accounts of wildfire dynamics or contextual details — which are critical for understanding evacuees’ perceptions and the rationale behind their evacuation decisions. To address these limitations, the recent emergence of Large Language Models (LLMs) provides exceptional reasoning capabilities to model and predict evacuation decision-making processes Huang and Chang (2022); Li et al. (2024e, d); Nguyen et al. (2024); Liu et al. (2024); Lee et al. (2024). Compared to traditional statistical models, LLMs display theory of mind (ToM) capabilities and have the potential to bridge the information gap present in survey data by better approximating human decision-making logic. LLMs also facilitate the integration of contextual information into the predictive process.
@@
However, employing the existing LLM framework for evacuation decision-making modeling and prediction, with socio-demographic information as inputs, still faces four significant challenges: (1) Mismatching with behavioral theory: Evacuation survey data size is often limited, for example, 334 valid examples for the 2021 Marshall wildfires Forrister et al. (2024a). LLMs tend to capture only partial reasoning patterns and overfit limited survey data, struggling to align with established behavioral theories Tjuatja et al. (2023); Petrov et al. (2024); Macmillan-Scott and Musolesi (2024). (2) Conflicting preferences in aligning with human thought: Reinforcement Learning with Human Feedback (RLHF) offers a promising approach to aligning LLM reasoning with human thought Sun et al. (2023); Zhang et al. (2024); Xu et al. (2024b). However, it is still challenging to accommodate individuals with diverse evacuation patterns Zhao et al. (2022); Sun et al. (2024b). For example, some individuals may evacuate immediately upon receiving an official order, prioritizing institutional guidance, while others may rely on social cues, choosing to stay until observing their neighbor evacuating. (3) Incorporating implicit mental states: Previous studies show that integrating mental states will benefit the improvement of human behavior predictions Gu et al. (2024). However, in a highly dynamic and chaotic wildfire environment, there exist many implicit, diverse, and complex mental states, perceptions, or beliefs, that drive wildfire evacuation behaviors, making it challenging to learn and incorporate such implicit mental states. (4) One mental state may map to multiple behaviors: Even if individuals share the same mental state, their behaviors can vary due to external factors. For instance, two individuals with identical risk perceptions might make different evacuation decisions due to the different traffic congestion situations and shelter accessibility Collins et al. (2018); Hong and Frias-Martinez (2020).
@@
The proposed framework adapts empirical psychological and behavioral knowledge to inform the variable selection, CoT template construction, and model inference to constrain the over-expressiveness of LLMs on small datasets and encourage a better alignment with existing behavioral theories. Both reasoning path classifier and memory-based RL help mitigate RLHF’s conflicting preference issues and tailor the prediction to individual behavior. Moreover, we calibrate implicit mental state generation chains through self-validation with a classifier based on PADM and sparsely available but implicit risk/threat-related answers in the survey data. We also augment the evacuation decision prediction with similar past error trials and their reflections from memory, guiding the model toward more accurate outcomes. We also integrate descriptive and descriptive external knowledge (e.g., wildfire progression) as external cues about the individual situation, guiding the LLM to translate the inferred mental state into an accurate behavior prediction. Our main contributions include:
  • We propose a novel framework that integrates LLMs’ advanced reasoning capability with psychological and behavioral theories, improving evacuation decision prediction accuracy in small, highly imbalanced datasets.
  • We introduce a behavioral theory-informed classifier to distinguish individual reasoning patterns, addressing conflicting preference problems while constraining the reasoning paradigm.
@@ 2 Related Work
2.1 Wildfire Evacuation Decision Prediction
Recent research has employed multiple methods to predict wildfire evacuation decisions. McCaffrey et al. (2018) employed a multinomial logistic model based on PADM, enhanced by a latent class approach, to predict diverse evacuation decisions in three U.S. fire-prone counties. Forrister et al. (2024a) applied logistic and linear regression to predict risk perception, evacuation decision, and delay time. Xu et al. (2023) benchmarked seven machine learning approaches (e.g., Random Forest, Classification And Regression Trees (CART), Extreme Gradient Boosting) and identified CART as the top-performing model for predicting evacuation behavior from the 2019 Kincade Fire survey. Meanwhile, Lovreglio et al. (2020) introduced the Wildfire Decision Model (WDM) calibrated via Hybrid Choice Models (HCM), incorporating latent factors like risk perception and prior experience for more accurate evacuation decision predictions. Sun et al. (2024c) further integrates risk perception and threat assessment as latent variables into an HCM framework, improving prediction accuracy. Traditional statistical models do not account for the logical flow of decision-making. HCM, in contrast, considers this process.
@@ 2.2 LLMs for Human Decision and Behavior Prediction
Current work increasingly leverages LLMs to model and predict human decisions. BigToM Gandhi et al. (2024) evaluates LLMs’ Theory-of-Mind (ToM) reasoning via causal templates and finds that GPT-4 partially matches human ToM performance while other models lag behind. Similarly, SUVA Leng and Yuan (2023) utilizes probabilistic modeling and behavioral economics games, revealing that larger LLMs display stronger prosocial and group-identity effects. DEBATunE Li et al. (2024a) utilizes a multi-agent debate process for data synthesis on controversial debate topics and utilizes supervised fine-tuning to simulate behaviors on controversial topics. Extending these insights, ToM Amirizaniani et al. (2024) focuses on open-ended social reasoning from Reddit’s ChangeMyView posts, showing prompt tuning with human intentions and emotions boosts performance but remains below human-level comprehension. Although these models can infer mental states, T4D Zhou et al. (2023) highlights the challenge of converting such inferences into strategic action, as even GPT-4 struggles without structured guidance. In parallel, LELMA Mensfelt et al. (2024) integrates symbolic AI to verify logical consistency in social simulations like the Prisoner’s Dilemma, demonstrating that self-refinement methods can improve the reliability of LLM-generated reasoning. Meanwhile, SimpleToM Gu et al. (2024) underscores that while LLMs can predict mental states and behavior, they often require deliberate prompting for accurate moral or behavioral judgments. Kang et al. (2023) propose the Value Injection Method (VIM), embedding human core values to enhance opinion and choice prediction. However, Kuribayashi et al. (2024) caution that instruction tuning and prompting do not inherently offer better alignment with human cognition than direct probability outputs from base LLMs. Zhu et al. (2024) reveal that arithmetic-trained LLMs can surpass classic decision-theoretic models when evaluating risky and time-delayed choices, demonstrating that specialized numerical training augments behavioral prediction. Liu et al. (2024) emphasizes a core limitation: LLMs systematically assume people behave more rationally than they do, underestimating well-documented human biases and highlighting a persistent gap between model predictions and real-world decision-making.
@@ 3.1 Preliminary
Wildfire Evacuation Survey Data: The Wildfire Evacuation Survey collects data about participants’ experiences during wildfire events through a set of multiple-choice, scaled, and open-ended questions. Questions Kuligowski et al. (2022a) cover topics such as prior wildfire risk knowledge, emergency plans, evacuation experiences, property damage, warning system awareness, and household socio-demographics. Responses range from numerical scales (e.g., personal injury perception rated from 1 to 5) and categorical choices (e.g., “Yes” or “No”) to counts (e.g., number of evacuations), while also capturing qualitative details like medical conditions and household composition. For detailed and complete survey refer to Kuligowski et al. (2022b) for Kincade Fire, and  Forrister et al. (2024b) for the Marshall Fire.
@@ 3.2 Reasoning Process Formulation
In this section, we describe how we construct classifiers based on PADM that identify the most probable reasoning patterns from survey data variables. These patterns are derived from the previously introduced risk perception and threat assessment. Once the appropriate reasoning pattern is determined, we demonstrate how an LLM generates the corresponding perception and integrates it into a CoT template, yielding the finalized CoT for evacuation prediction.
@@ 3.2.1 Variable Selection for Perceptions
Building on the HCM framework based on PADM for wildfire evacuation decision prediction Lovreglio et al. (2020); Sun et al. (2024c), our approach aims to develop a unified statistical method to automatically select the key variables that contribute to risk perceptions and threat assessment by examining all available survey questions.
As briefly mentioned in Section 3.1, the survey includes questions capturing socio-demographic data, awareness and understanding, and decision-related factors such as prior wildfire risk awareness, emergency preparedness, evacuation experience, warning system awareness, personal injury perceptions, household income, employment status, and medical conditions. In the HCM method Sun et al. (2024c), threat assessment and risk perceptions are validated using specific indicators derived from survey questions, requiring manual selection of variables, fitting them to indicators, and evaluating their alignment. Our approach automates this process by incorporating all available variables into the fitting process and selecting those with the highest weights, ensuring the strongest contributions to risk perceptions and threat assessments. This minimizes bias in manual selection and enhances the model’s ability to capture key evacuation decision factors.
@@
Empirically, $\theta$ corresponds to the elbow point in the weight distribution, ensuring key variables are retained while filtering out less significant ones. The empirical results visualizing this elbow point are provided in Appendix E, demonstrating the sharp decline in variable importance beyond the selected threshold.
@@ 3.2.2 CoT Construction based on Perceptions
In this section, we describe how we construct the CoT by leveraging the variable subsets selected earlier. First, we develop a universal CoT template (see Appendix B) that structures reasoning into two stages — threat assessment followed by risk perception — to enforce a behaviorally grounded reasoning path. We then introduce a reasoning pattern classifier that selects the most probable reasoning pattern for each individual based on the prediction success rate of each reasoning pattern. Finally, we prompt the LLM to generate textual threat assessments and risk perceptions along with their quantitative perception scores for each case, calibrating consistency and establishing a foundation for subsequent evacuation decision prediction.
@@
Perception Inference: After selecting each individual’s most probable reasoning pattern, we refine the generation of risk perceptions and threat assessments to construct a complete CoT for evacuation decision prediction. The LLM first generates textual threat assessment and risk perceptions while explicitly assigning quantitative perception scores (1–5) as calibration indicators. This dual representation enhances consistency between inferred perceptions and key survey variables. To refine this calibration, we use the first 70% of the dataset to establish a knowledge base mapping LLM-generated perceptions to survey-derived scores, allowing the model to align qualitative reasoning with numerical assessment.
@@
Once robust textual perceptions and their corresponding numerical scores are obtained, we integrate them into the CoT templates along with external information such as environment context derived from survey questions (e.g., “I’m not in the area ordered to evacuate.”) and user inputs (detailed in Figure 4), forming the complete CoT for evacuation decision prediction. This construction process ensures that the CoT consistently adapts to each individual’s most probable reasoning pattern while leveraging textual information.
With this complete CoT in place, we establish a reasoning process that accommodates multiple threat–risk scenarios, setting the stage for aligning the reasoning process more closely with human decision-making through the memory-based RL module.

### v2

@@ Abstract
Ruxiao Chen*1, Chenguang Wang1thanks: Equal Contribution., Yuran Sun2, Xilei Zhao2, Susu Xu1 1Johns Hopkins University 2University of Florida {rchen117, cwang274, susuxu}@jhu.edu
@@ 1 Introduction
Wildfires are emerging as a significant natural hazard worldwide Jain et al. (2020); Zahura et al. (2024). In the January 2025 Southern California wildfires, more than 200,000 residents received evacuation orders to leave their homes Stelloh et al. (2025). There is an urgent demand for emergency planners and policymakers to develop effective evacuation strategies to mitigate wildfire impacts Mockrin et al. (2018); Tapley et al. (2023). However, successful evacuations require a clear understanding of the human evacuation decision-making process and outcomes (i.e. whether individuals will follow the order to evacuate or stay) during these events to help policymakers improve evacuation order design, develop more efficient emergency response strategies, and build more resilient communities Collins et al. (2018); Lovreglio et al. (2020); Hong and Frias-Martinez (2020); Sun et al. (2024c). risksLovreglio et al. (2020); Sun et al. (2024c).
@@
Previous studies often construct evacuation choice models through a conceptual framework, Protective Action Decision Model (PADM) Strahan and Watson (2019); Lovreglio et al. (2019); Santana et al. (2021); Sun et al. (2024a), which are designed to incorporate psychological factors, like individual risk perception and threat assessment, into the prediction process. Based on the PADM framework, past methods may employ various statistical models (e.g., logistic regression Forrister et al. (2024a), multinomial logistic regression McCaffrey et al. (2018)) to predict individual-evacuation decisions using socio-demographic information as inputs, trained on post-wildfire survey data. However, these traditional PADM-type statistical models lack reasoning capabilities to capture the diverse and complex logic underlying human decision-making due to limited data and restrictive modeling structure, even when the survey design is grounded in established behavioral theories. In addition, these statistical methods struggle to integrate qualitative descriptions, such as narrative accounts of wildfire dynamics or contextual details, which are critical for understanding evacuees’ perceptions and the rationale behind their evacuation decisions.
To address these limitations, the recent emergence of Large Language Models (LLMs) provides exceptional reasoning capabilities to model and predict evacuation decision-making processes Huang and Chang (2022); Nguyen et al. (2024); Liu et al. (2024); Lee et al. (2024). Compared to traditional statistical models, LLMs display theory of mind (ToM) capabilities and have the potential to bridge the information gap present in survey data by better approximating human decision-making logic. LLMs also facilitate the integration of contextual information into the predictive process.
However, employing the existing LLM framework for evacuation decision-making modeling and prediction, with socio-demographic information as inputs, still faces four significant challenges: (1) Mismatching with behavioral theory: Evacuation survey data size is often limited, for example, 334 valid examples for the 2021 Marshall wildfires Forrister et al. (2024a). LLMs tend to capture only partial reasoning patterns and overfit limited survey data, struggling to align with established behavioral theories Tjuatja et al. (2023); Petrov et al. (2024); Macmillan-Scott and Musolesi (2024).
@@
(2) Conflicting preferences in aligning with human thought: Reinforcement Learning with Human Feedback (RLHF) offers a promising approach to aligning LLM reasoning with human thought Sun et al. (2023); Zhang et al. (2024); Xu et al. (2024b). However, it is still challenging to accommodate individuals with diverse evacuation patterns Zhao et al. (2022); Sun et al. (2024b). For example, some individuals may evacuate immediately upon receiving an official order, prioritizing institutional guidance, while others may rely on social cues, choosing to stay until observing their neighbor evacuating. (3) Incorporating implicit mental states: Previous studies show that integrating mental states will benefit the improvement of human behavior predictions Gu et al. (2024). However, in a highly dynamic and chaotic wildfire environment, there exist many implicit, diverse, and complex mental states, perceptions, or beliefs, that drive wildfire evacuation behaviors, making it challenging to learn and incorporate such implicit mental states. (4) One mental state may map to multiple behaviors: Even if individuals share the same mental state, their behaviors can vary due to external factors. For instance, two individuals with identical risk perceptions might make different evacuation decisions due to the different traffic congestion situations and shelter accessibility Collins et al. (2018); Hong and Frias-Martinez (2020).
@@
The proposed framework adapts empirical psychological and behavioral knowledge to inform the variable selection, CoT template construction, and model inference to constrain the over-expressiveness of LLMs on small datasets and encourage a better alignment with existing behavioral theories. Both reasoning path classifier and memory-based RL help mitigate RLHF’s conflicting preference issues and tailor the prediction to individual behavior. Moreover, we calibrate implicit mental state generation chains through self-validation with a classifier based on PADM and sparsely available but implicit risk/threat-related answers in the survey data. We also augment the evacuation decision prediction with similar past error trials and their reflections from memory, guiding the model toward more accurate outcomes. We also integrate descriptive external knowledge (e.g., wildfire progression) as external cues about the individual situation, guiding the LLM to translate the inferred mental state into an accurate behavior prediction.
Our main contributions include:
  • We introduce a novel framework that integrates advanced reasoning capability of LLMs with psychological and behavioral theories, improving the accuracy of evacuation decision prediction in small, highly imbalanced data sets.
  • We design a behavioral theory-informed classifier to distinguish individual reasoning patterns, addressing conflicting preference problems while constraining the reasoning paradigm.
@@ 2 Related Work
2.1 Disaster Evacuation Decision Prediction
Sudden-onset natural hazards—such as wildfires, hurricanes, and earthquakes, often trigger cascading failures that result in widespread environmental damage and human displacement Xu et al. (2022); Li et al. (2023b); Wang et al. (2024b). In response, critical tasks such as hazard progression tracking Chen et al. (2024b), damage assessment Yu et al. (2024); Xue et al. (2024); Li et al. (2025); Li and Xu (2025), and large-scale evacuation become urgent priorities to mitigate impacts and minimize losses. Recent research has employed multiple methods to predict wildfire evacuation decisions. McCaffrey et al. (2018) employed a multinomial logistic model based on PADM, enhanced by a latent class approach, to predict various evacuation decisions in three US fire-prone counties. Forrister et al. (2024a) applied logistic and linear regression to predict risk perception, evacuation decision, and delay time. Xu et al. (2023) benchmarked seven machine learning approaches (e.g., Random Forest, Classification And Regression Trees (CART), Extreme Gradient Boosting) and identified CART as the best-performing model for predicting evacuation behavior from the 2019 Kincade Fire survey. Meanwhile, Lovreglio et al. (2020) introduced the Wildfire Decision Model (WDM) calibrated via Hybrid Choice Models (HCM), incorporating latent factors like risk perception and prior experience for more accurate evacuation decision predictions. Sun et al. (2024c) further integrates risk perception and threat assessment as latent variables into an HCM framework, improving prediction accuracy. Traditional statistical models do not account for the logical flow of decision-making. HCM, in contrast, considers this process.
@@ 2.2 LLMs for Human Decision and Behavior Prediction
Recent work increasingly leverages LLMs to model and predict human decision-making. BigToM Gandhi et al. (2024) evaluates LLMs’ Theory-of-Mind (ToM) capabilities using causal templates, finding that GPT-4 partially approximates human ToM reasoning, while other models lag behind. SUVA Leng and Yuan (2023) applies probabilistic modeling to behavioral economics games, revealing that larger LLMs exhibit stronger prosocial and group-identity effects. SUVA Leng and Yuan (2023) and  Amirizaniani et al. (2024) find that larger LLMs capture prosocial behavior and emotional reasoning, though still fall short of human-level comprehension. T4D Zhou et al. (2023) highlights LLMs’ difficulty in translating inferred mental states into strategic action without structured guidance. LELMA Mensfelt et al. (2024) improves reasoning reliability through symbolic consistency checks, while SimpleToM Gu et al. (2024) emphasizes the need for deliberate prompting to elicit accurate moral and behavioral judgments. SimpleToM Gu et al. (2024) shows that while LLMs can predict mental states and behavior, they often require careful prompting to yield accurate moral or behavioral judgments. To address this, Kang et al. (2023) propose the Value Injection Method (VIM), embedding core human values into model outputs. However, Kuribayashi et al. (2024) argue that prompting does not inherently yield better cognitive alignment than base model probabilities. Zhu et al. (2024) find that arithmetic-trained LLMs can outperform classical models in evaluating risky, time-delayed choices. Still, Liu et al. (2024) point out that LLMs tend to overestimate human rationality, overlooking well-documented cognitive biases and limiting behavioral fidelity.
A recent survey paper Lei et al. (2025) reviews a growing body of work exploring the use of LLMs in disaster management, utilizing methods such as retrieval-augmented generation (RAG), instruction tuning, prompt chaining, and knowledge graph grounding. These approaches support tasks such as classifying social media content, estimating human loss Wang et al. (2024a), generating summaries, and answering public queries, thus improving situational awareness and response coordination. Specifically, Chen et al. (2024a) proposed E-KELL, a framework that grounds large language models in structured knowledge graphs constructed from disaster domain knowledge. This design improves the reliability and interpretability of disaster response decisions by embedding this rule-based knowledge into the LLM reasoning process. Yin et al. (2025) introduced CrisisSense-LLM, which instruction-tunes LLaMA2 on a large corpus of disaster-related tweets to perform multi-label classification over event type, informativeness, and aid-related categories, thereby improving situational awareness in disaster contexts.
While these studies demonstrate the growing utility of LLMs in disaster informatics, they primarily focus on information extraction, structured representation, and retrieval. In contrast, our work centers on modeling human behavior under disaster stress by integrating behavioral theory and cognitive perception constructs into the LLM prompting process—enabling the simulation of individual decision-making pathways such as evacuation under threat.
@@ 3.1 Preliminary
Wildfire Evacuation Survey Data: The Wildfire Evacuation Survey collects data about participants’ experiences during wildfire events through a set of multiple-choice, scaled and open-ended questions. The questions Kuligowski et al. (2022a) cover topics such as prior knowledge of wildfire risk, emergency plans, evacuation experiences, property damage, warning system awareness, and household socio-demographics. Responses range from numerical scales (e.g., perception of personal injury rated to 5) and categorical choices (e.g. Yes” or “No”) to counts (e.g., number of evacuations), while also capturing qualitative details like medical conditions and household composition. For detailed and complete survey refer to Kuligowski et al. (2022b) for Kincade Fire, and  Forrister et al. (2024b) for the Marshall Fire.
@@ 3.2 Reasoning Process Formulation
In this section, we describe how we construct classifiers based on PADM that identify the most probable reasoning patterns from survey data variables. These patterns are derived from the previously introduced risk perception and threat assessment. Once a reasoning pattern is identified, we demonstrate how an LLM generates the corresponding perception and integrates it into a CoT template, yielding the finalized CoT for evacuation prediction.
@@ 3.2.1 Variable Selection for Perceptions
Building on the HCM framework grounded in PADM for wildfire evacuation decision prediction Lovreglio et al. (2020); Sun et al. (2024c), our approach aims to develop a unified statistical method to automatically select the key variables that contribute to risk perceptions and threat assessment by examining all available survey questions.
As briefly mentioned in Section 3.1, the survey includes questions capturing socio-demographic data, awareness and understanding, and decision-related factors such as prior wildfire risk awareness, emergency preparedness, evacuation experience, warning system awareness, personal injury perceptions, household income, employment status, and medical conditions.
In the original HCM method Sun et al. (2024c), threat and risk perceptions are validated using specific indicators derived from survey questions, requiring manual selection of variables, fitting them to indicators, and evaluating their alignment. Our approach automates this process by incorporating all available variables into the fitting process and selecting those with the highest weights, ensuring the strongest contributions to risk perceptions and threat assessments. This minimizes bias in manual selection and enhances the model’s ability to capture key evacuation decision factors.
@@
Empirically, $\theta$ corresponds to the elbow point in the weight distribution, ensuring key variables are retained while filtering out less significant ones. The empirical results visualizing this elbow point are provided in Appendix E, demonstrating the sharp decline in variable importance beyond the selected threshold. To justify the necessity of selecting representative variable subsets for each perception, we conducted an ablation experiment using all survey variables without filtering. This resulted in a notable decline in performance, attributable to the transformer’s limited capacity to maintain focused attention over long and densely encoded inputs. See Appendix C (Table 9) for further details.
@@ 3.2.2 CoT Construction based on Perceptions
This section outlines our approach to constructing the Chain-of-Thought (CoT) using the variable subsets identified earlier. We first develop a universal CoT template (Appendix B) that organizes reasoning into two behaviorally grounded stages: threat assessment followed by risk perception. Next, we introduce a reasoning pattern classifier that selects the most likely reasoning path for each individual based on prediction success rates across candidate patterns. Finally, we prompt the LLM to generate textual threat and risk perceptions alongside quantitative scores, establishing consistency and providing the foundation for downstream evacuation decision prediction.
@@
Perception Inference: After selecting each individual’s most probable reasoning pattern, we prompt the LLM to generate the corresponding threat perceptions and risk perceptions to construct a complete CoT for evacuation decision prediction. The LLM first generates textual threat and risk perceptions while explicitly assigning quantitative perception scores (1–5) as calibration indicators. This dual representation enhances consistency between inferred perceptions and key survey variables. To refine this calibration, we use the first 70% of the dataset to build a knowledge base that maps LLM-generated perceptions to survey-derived scores.
@@
The resulting textual perceptions and calibrated scores are then integrated into the CoT template, along with contextual information extracted from survey responses (e.g., “I’m not in the area ordered to evacuate”) and user-specific inputs (Table 7), forming an individualized and perception-grounded reasoning chain for decision modeling.